verify that each given function is a solution of the

differential equation.

9.t^2y′′\:+\:5ty′\:+\:4y\:=\:0;\:y\left(t\right)\:=\:t^{−2}


1
Expert's answer
2021-09-05T18:13:42-0400

Let us verify that the function "\:y\left(t\right)\:=\:t^{−2}" is a solution of the differential equation "t^2y''\:+\:5ty'\:+\:4y\:=\:0." Taking into account that "y'(t)=-2t^{-3}, \ y''(t)=6t^{-4}," and

"t^2y''(t)+\:5ty'(t)+4y=t^2(6t^{-4})+\:5t(-2t^{-3})+4t^{-2}=6t^{-2}-10t^{-2}+4t^{-2}=0,"

we conclude that the function "\:y\left(t\right)\:=\:t^{−2}" is a solution of the differential equation "t^2y''\:+\:5ty'\:+\:4y\:=\:0."